Find the odds against obtaining two (2) heads in a single toss of two coins.
Is it 1/2

Found 2 solutions by stanbon, venugopalramana:
Answer by stanbon(48568)

Find the odds against obtaining two (2) heads in a single toss of two coins.
"Odds against event A" = [P(not A)]/[P(A)]
P(2 heads) = 1/4
P( not 2 heads) = 3/4
Therefore Odds against 2 heads = (3/4)/(1/4) = 3:1
Cheers,
Stan H.

Answer by venugopalramana(3286)

Find the odds against obtaining two (2) heads in a single toss of two coins.
Is it 1/2...NO...
THERE ARE 2 POSSIBILITIES IN EACH TOSS ...PROBABILITY OF HEAD IN ONE TOSS=1/2
SAME IS CASE IN SECOND TOSS =1/2
HENCE PROBABILITY OF GETTING 2 HEADS IN 2 TOSSES =(1/2)*(1/2)=1/4
PROBABILITY AGAINST 2 HEADS =1-1/4=3/4
ODDS AGAINST 2 HEADS ARE 3/4:1/4=3:1